OpenStudy (anonymous):

which expression best describes these equations? A. a(b + c) = ab + ac B. This is what the 5 in the term 5xy is called. C. (a + b) + c = a + (b + c) D. This expression has 4 of them: 3a + b - 5 + 8d. E. a + b = b + a

OpenStudy (anonymous):

A. a(b + c) = ab + ac ( distributive property ) ----------- B. This is what the 5 in the term 5xy is called. (coefficient)----- C. (a + b) + c = a + (b + c) ( associative property ) ----- D. This expression has 4 of them: 3a + b - 5 + 8d. (terms)------ E. a + b = b + a (commutative prperty)